# On Steerability Factors for Growing Vine Robots

### v1

@@ Abstract
Vine robots extend their tubular bodies by everting material from the tip, enabling navigation in complex environments with a minimalist soft body. Despite their promise for field applications, especially in the urban search and rescue domain, performance is constrained by the weight of attached sensors or tools, as well as other design and control choices. This work investigates how tip load, pressure, length, diameter, and fabrication method shape vine robot steerability—the ability to maneuver with controlled curvature–for robots that steer with series pouch motor-style pneumatic actuators. We conduct two groups of experiments: (1) studying tip load, chamber pressure, length, and diameter in a robot supporting itself against gravity, and (2) studying fabrication method and ratio of actuator to chamber pressure in a robot supported on the ground. Results show that steerability decreases with increasing tip load, is best at moderate chamber pressure, increases with length, and is largely unaffected by diameter. Robots with actuators attached on their exterior begin curving at low pressure ratios, but curvature saturates at high pressure ratios; those with actuators integrated into the robot body require higher pressure ratios to begin curving but achieve higher curvature overall. We demonstrate that robots optimized with these principles outperform those with ad hoc parameters in a mobility task that involves maximizing upward and horizontal curvatures.
@@ I Introduction
Robots are routinely used to gather important information from dangerous environments; to be effective, they must move through these environments while carrying the sensors and tools needed for their task. Vine robots, soft inflatable robots that extend by everting their body from their tip [1], have shown promise in navigating confined spaces. They have previously explored urban search and rescue training sites [2, 3], an archaeological site [4], and a salamander burrow habitat [5]. They are also capable of squeezing through small apertures [6] and growing vertically [4]. However, these demonstrations largely relied on ad hoc design and control parameters, limiting their generalization.
Vine robot mobility is constrained when they carry tip-mounted loads [3] or must span open spaces [7]. Because tip mounts, sensors, and tools add mass while altering structural and control demands, it is essential to systematically quantify how design and control parameters—such as tip weight, internal pressures, length, diameter, and body fabrication—influence steerability. Steerability, defined here as the ability of a vine robot to generate controlled curvature in free space, is critical for maneuvering in complex environments and for ensuring reliable field performance.
@@
This work addresses this gap by experimentally characterizing how vine robot steerability depends on design and control parameters (Fig. 1). Specifically, we focus on vine robots that steer using series pouch motor-style pneumatic actuation [4, 8]. We examine (1) how tip load, chamber pressure, length, and diameter affect a robot supporting itself against gravity; and (2) how actuator placement (exterior [7] vs. integrated [8]) and the ratio of actuator to chamber pressure affect planar curvature in robots supported on the ground. While prior studies have characterized unweighted curvature [9], growth speed [10], and steering force [11] for vine robots, our focus is identifying trends in 3D steerability under loads, enabling the development of generalizable design and control principles for pneumatically steered vine robots.
@@ II Related Work
Continuum robot research has shown that where and how actuation is applied strongly influences deformation and workspace [12, 13]. For vine robots, the choice of steering mechanism has a significant impact on the curvature the robot can achieve. Discrete methods such as latches [6] or locally placed magnets [14] provide sharp curvatures at predefined locations, but are challenging to reverse. Internal tip devices [2, 15] allow high-curvature steering of only the robot tip, which is less useful for navigating free space. External magnets [16] can steer vine robots, but only in small-scale environments where magnetic forces can reach.
Various pneumatic actuators offer reversible, continuous curvature steering for vine robots, including series pneumatic artificial muscles [17], fabric pneumatic artificial muscles [18], and cylindrical pneumatic artificial muscles [19]. Series pouch motors [4, 8] are another type of pneumatic actuator that are easily fabricated using impulse heat-sealing or heat pressing [20]. These actuators can be interfaced with the main vine robot main chamber tube either as external elements attached to the body wall [4] or integrated within the wall itself [8]. The quantitative impact of fabrication methods on free-space steerability has not been studied.
Other parameters relevant to free-space steerability include tip load, internal pressure, and body geometry. Vine robots must balance the need to carry sensors and tools at the tip with the risk of collapse under transverse loading or axial buckling. Increasing a vine robot’s internal pressure or diameter [7], or incorporating stiffening elements [21] helps it better support its own weight but also affects its ability to steer. Previous studies have looked at how vine robots steer when using obstacles to guide their path on flat ground [22, 23] and have compared curvatures that can be achieved with various pneumatic actuators [9] without tip loading. No study has systematically compared design and control parameters for 3D mobility tasks in free space for series pouch motor-actuated robots. We expect our results to directly inform other pneumatically steered vine robots, and to provide helpful insights for vine robots steered through other means.
@@ III-B Experimental Setup
Figure 2: Experimental setup and points of interest for self-supporting 3D steerability experiments. (a) The robot is placed on a table with support only underneath the base section. Tracking the base LED, tip LED, and coordinate axis allows us to align the data with a reference frame fixed to the table. (b) The order in which the robot tip visits the six via points is shown in blue. The convex hull area of those points is shown in red. The square root of the convex hull area is the characteristic length, which is plotted in Fig. 3, along with the horizontal and vertical range.
Figure 3: Results of self-supporting 3D steerability experiments. For all parameters, we plot the average characteristic length and its standard deviation, along with the horizontal and vertical range for each trial. (a) Increasing tip load significantly decreases characteristic length, horizontal range, and vertical range. (b) Increasing pressure initially increases and then decreases characteristic length; it significantly decreases horizontal range and changes vertical range erratically. (c) Increasing robot length increases characteristic length and significantly increases horizontal range, while decreasing vertical range. (d) As long as the robot is not collapsed (leftmost data point), varying diameter does not significantly impact characteristic length, horizontal range, or vertical range.
@@
For each trial, the robot moved through six via points (red dots connected by a blue path in Fig. 2(b)). We inflated each actuator in turn to its maximum pressure of 10.34 kPa (determined using a safe margin from the pressure at which we expected the heat seals to tear), as well as each pairing of two actuators to their maximum pressure at the same time, with the other actuators set to zero pressure. We used four closed-loop pressure regulators (QB3, Proportion-Air, McCordsville, IN) to enforce the desired values for chamber and actuator pressures. Before commanding the next set of pressures, we waited for the robot to stop moving. We conducted three trials for each set of parameters. The chamber was deflated between trials, and the robot was reset to have all actuators at zero pressure, to ensure a more repeatable path.
We used a motion capture (Impulse X2E, PhaseSpace, San Leandro, CA) LED marker to track the 3D tip position of the robot. We also used LEDs aligned with our desired coordinate frame and an LED at the base of the robot to realign the data to our global reference frame located at the base of the robot with the $z$-axis pointing along the robot toward its tip and the $y$-axis pointing up.
To visualize the steerability of the robot, we project the 3D tip position data into the $x$-$y$ plane and calculate the convex hull area of the region enclosed by the $x$ and $y$ coordinates of $m$ marker positions (red shape in Fig. 2(b)) using Gauss’s area formula. We then take the square root of the convex hull area to yield characteristic length:
$L_{c}=\sqrt{\tfrac{1}{2}\left|\sum_{k=1}^{m}(x_{k}y_{k+1}-x_{k+1}y_{k})\right|},$ (1)
where $x_{m+1}=x_{1}$ and $y_{m+1}=y_{1}$. We plot the average characteristic length across three trials and its standard deviation, along with the horizontal range (the difference between the minimum and maximum $x$ values) and the vertical range (the difference between the minimum and maximum $y$ values) for each trial (Fig. 3).
@@ III-C Tip Load
When varying tip load, we maintained the following constant parameters: chamber pressure of 6.89 kPa: high enough to support the robot while also being low enough that it steers freely; length of 0.61 m: sufficiently long to steer without causing the robot to excessively deflect under its own weight; and inflated diameter of 8.1 cm: close to the size of previously deployed field robots [4, 2, 7]. The robot performed its task under loads of 0 g, 20 g, 50 g, 100 g, 150 g, and 200 g (the maximum mass the robot could hold before it started approaching collapse behavior).
Steerability as a function of tip load is shown in Fig. 3(a). There is a clear decrease in characteristic length with increased tip load, with the drop becoming especially significant at 100 g and above. Correspondingly, the horizontal and vertical range experience sharp decreases as the tip load increases.
@@ III-D Chamber Pressure
When varying chamber pressure, we maintained the following constant parameters: tip load of 50 g: an intermediate load that causes some deflection but not so little steerability as to prevent analysis of other parameters; length of 0.61 m; and diameter of 8.1 cm. The robot performed its task at chamber pressures of 2.07 kPa (close to the minimum chamber pressure that supports the robot), 3.45 kPa, 5.52 kPa, 6.89 kPa, 8.27 kPa, and 10.34 kPa (close to the maximum chamber pressure it holds without bursting).
Steerability as a function of chamber pressure is shown in Fig. 3(b). The characteristic length starts high at a low pressure and initially increases with pressure, peaking at 5.52 kPa. Beyond this point, higher pressures lead to a decreased characteristic length, as the robot is too stiff for the actuators to move it effectively. The horizontal range generally decreases with increasing pressure, which is expected with the robot’s increasing stiffness. The vertical range generally increases with increasing pressure but experiences high variation. This is likely due to the fact that the robot is becoming stiffer and harder to move, but that same stiffness is also preventing deflection and therefore allowing the robot to make larger motions if it creates enough force to do so.
@@ III-E Length
When varying robot length, we maintained the following constant parameters: tip load of 50 g; chamber pressure of 6.89 kPa; and diameter of 8.1 cm. The robot performed its task at lengths of 0.30 m, 0.46 m, 0.61 m, and 0.76 m.
Steerability as a function of robot length is shown in Fig. 3(c). Generally, the characteristic length increases with increasing length. There is also a clear increase in horizontal range, which makes sense, since a longer robot’s tip should travel farther for a given amount of curvature. However, the decreasing vertical range shows that this increase in length results in a heavier robot that struggles more to move against gravity. Eventually, the robot would become so long that it would collapse, reducing the characteristic length, horizontal range, and vertical range to nearly zero.
@@ III-F Diameter
When varying diameter, we maintained the following constant parameters: tip load of 50 g; chamber pressure of 6.89 kPa; and length of 0.61 m. The robot performed its task at inflated diameters of 3.2 cm (smaller than the typical field exploration robot), 6.5 cm, 9.7 cm, and 12.9 cm (larger than the typical field exploration robot).
Steerability as a function of robot diameter is shown in Fig. 3(d). There is minimal change in characteristic length across diameters, aside from the 3.2 cm robot, which collapsed under the given parameters. Thus, having a diameter large enough to avoid collapse is important, but how much larger is of minimal consequence. Similarly, the horizontal range and vertical range are fairly constant once the robot is at a sufficient diameter to not collapse, though there is a slight decrease in horizontal range and an increase in vertical range with increasing diameter.
@@ IV-A Robot Body Fabrication
In the exterior pouch design [4], the main chamber is first fabricated as a single-layer thermoplastic polyurethane (TPU)-coated ripstop nylon tube. Three series pouch motor actuators, each with half the lay-flat width of the main chamber tube, and with pouches of equal length and width (experimentally determined to yield maximum contraction [9]) are then constructed and taped outside the main chamber.
In the integrated pouch design [8], the main chamber and actuators are built together as a double-layer structure. A folded piece of TPU-coated fabric is heat-sealed to define both the body wall and internal pouch chambers before being joined lengthwise into a tube. In this construction, each series pouch motor actuator has a nominal width of two-thirds of the lay-flat diameter of the main chamber tube.

### v2

@@ Abstract
Vine robots extend their tubular bodies by everting material from the tip, enabling navigation in complex environments. Despite their promise for field applications, performance is constrained by the weight of attached sensors, and design and control choices. This work investigates how tip load, pressure, length, diameter, and fabrication method shape vine robot steerability–the ability to maneuver with controlled curvature–for robots that steer with series pouch motor-style pneumatic actuators. We conduct two groups of experiments: (1) studying tip load, chamber pressure, length, and diameter in a robot supporting itself against gravity, and (2) studying fabrication method and ratio of actuator to chamber pressure in a robot supported on the ground. Results show that steerability decreases with increasing tip load, is best at moderate chamber pressure, increases with length, and is largely unaffected by diameter. Robots with actuators attached on their exterior begin curving at low pressure ratios, but curvature saturates at high pressure ratios; those with actuators integrated into the robot body require higher pressure ratios to begin curving but achieve higher curvature overall. We demonstrate that robots optimized with these principles outperform those with ad hoc parameters in a mobility task that involves maximizing upward and horizontal curvatures.
@@ I Introduction
Robots are routinely used to gather important information from dangerous environments; to be effective, they must move through these environments while carrying the sensors and tools needed for their task. Vine robots, soft inflatable robots that extend by everting their body from their tip [blumenschein2020design], have shown promise in navigating confined spaces. They have previously explored urban search and rescue training sites [der2021roboa, mcfarlandSPROUT], an archaeological site [CoadRAM2020], and a salamander burrow habitat [qin20253d]. They are also capable of squeezing through small apertures [HawkesScienceRobotics2017] and growing vertically [CoadRAM2020]. However, these demonstrations largely relied on ad hoc design and control parameters, limiting their generalization.
Vine robot mobility is constrained when they carry tip-mounted loads [mcfarlandSPROUT] or must span open spaces [mcfarland2023collapse]. Because tip mounts, sensors, and tools add mass while altering structural and control demands, it is essential to systematically quantify how design and control parameters—such as tip weight, internal pressures, length, diameter, and body fabrication—influence steerability. Steerability, defined here as the ability of a vine robot to generate controlled curvature in free space, is critical for maneuvering in complex environments and for ensuring reliable field performance.
@@
This work addresses this gap by experimentally characterizing how vine robot steerability depends on design and control parameters (Fig. 1). Specifically, we focus on vine robots that steer using series pouch motor-style pneumatic actuation [CoadRAM2020, Ataka2020ModelBased]. We examine (1) how tip load, chamber pressure, length, and diameter affect a robot supporting itself against gravity; and (2) how actuator placement (exterior [mcfarland2023collapse] vs. integrated [Ataka2020ModelBased]) and the ratio of actuator to chamber pressure affect planar curvature in robots supported on the ground. While prior studies have characterized unweighted curvature [kubler2024comparison], growth speed [alvarez2025high], and steering force [Mendoza2024HighCurvature] for vine robots, our focus is identifying trends in 3D steerability under loads, enabling the development of generalizable design and control principles for pneumatically steered vine robots.
@@ II Related Work
Continuum robot research has shown that where and how actuation is applied strongly influences deformation and workspace [russo2023continuum, seleem2023recent]. For vine robots, the choice of steering mechanism has a significant impact on the curvature the robot can achieve. Discrete methods such as latches [HawkesScienceRobotics2017] or locally placed magnets [li2021bioinspired] provide sharp curvatures at predefined locations, but are challenging to reverse. Internal tip devices [der2021roboa, lee2023high] allow high-curvature steering of only the robot tip, which is less useful for navigating free space. External magnets [kim2025external] can steer vine robots, but only in small-scale environments where magnetic forces can reach.
Various pneumatic actuators offer reversible, continuous curvature steering for vine robots, including series pneumatic artificial muscles [greer2017series], fabric pneumatic artificial muscles [Naclerio2020simple], and cylindrical pneumatic artificial muscles [kubler2023multi]. Series pouch motors [CoadRAM2020, Ataka2020ModelBased] are another type of pneumatic actuator that are easily fabricated using impulse heat-sealing or heat pressing [mack2025efficient]. These actuators can be interfaced with the main vine robot main chamber tube either as external elements attached to the body wall [CoadRAM2020] or integrated within the wall itself [Ataka2020ModelBased]. The quantitative impact of fabrication methods on free-space steerability has not been studied.
Other parameters relevant to free-space steerability include tip load, internal pressure, and body geometry. Vine robots must balance the need to carry sensors and tools at the tip with the risk of collapse under transverse loading or axial buckling. Increasing a vine robot’s internal pressure or diameter [mcfarland2023collapse], or incorporating stiffening elements [do2024stiffness] helps it better support its own weight but also affects its ability to steer. Previous studies have looked at how vine robots steer when using obstacles to guide their path on flat ground [GreerICRA2018, selvaggio2020obstacle] and have compared curvatures that can be achieved with various pneumatic actuators [kubler2024comparison] without tip loading. No study has systematically compared design and control parameters for 3D mobility tasks in free space for series pouch motor-actuated robots. We expect our results to directly inform other pneumatically steered vine robots, and to provide helpful insights for vine robots steered through other means.
@@ III-B Experimental Setup
Figure 2: Experimental setup and points of interest for self-supporting 3D steerability experiments. (a) The robot is placed on a table with support only underneath the base section. Tracking the base LED, tip LED, and coordinate axis allows us to align the data with a reference frame fixed to the base of the robot. (b) The order in which the robot tip visits the six via points is shown in blue. The convex hull area of those points is shown in red.
Figure 3: Results of self-supporting 3D steerability experiments. For all parameters, we plot the paths traced out by the robot tip during each trial, and the average and standard deviation for the characteristic area, horizontal range, vertical range, and initial vertical position. (a) Increasing tip load significantly decreases characteristic length, horizontal range, vertical range, and initial vertical position. (b) Increasing pressure initially increases and then decreases characteristic length; it significantly decreases horizontal range and changes vertical range erratically. The initial vertical position generally increases. (c) Increasing robot length increases characteristic length and significantly increases horizontal range, while decreasing vertical range and initial vertical position. (d) As long as the robot is not collapsed (leftmost data point), varying diameter does not significantly impact characteristic length, horizontal range, vertical range, or initial vertical position.
@@
For each trial, the robot moved through six via points (red dots connected by a blue path in Fig. 2(b)), chosen to represent the full range of motion of the robot in the two degrees of freedom of its 3D tip position that are controlled by the steering actuators. We inflated each actuator in turn to its maximum pressure of 10.34 kPa (determined using a safe margin from the pressure at which we expected the heat seals to tear), as well as each pairing of two actuators to their maximum pressure at the same time, with the other actuators set to zero pressure. We used four closed-loop pressure regulators (QB3, Proportion-Air, McCordsville, IN) to enforce the desired values for chamber and actuator pressures. Before commanding the next set of pressures, we waited for the robot to stop moving. We conducted three trials for each set of parameters. The chamber was deflated between trials, and the robot was reset to have all actuators at zero pressure, to ensure a more repeatable path.
We used a motion capture (Impulse X2E, PhaseSpace, San Leandro, CA) LED marker to track the 3D tip position of the robot at a rate of 30 Hz. This tip-mounted LED is attached to the top of any weights on the robot on the top edge of the robot. We also used LEDs aligned with our desired coordinate frame and an LED on the top edge of the base of the robot to realign the data to our global reference frame located at the base of the robot with the $z$-axis pointing along the robot toward its tip and the $y$-axis pointing up.
To visualize the steerability of the robot, we project the 3D tip position data into the $x$-$y$ plane and calculate the convex hull area of the region enclosed by the $x$ and $y$ coordinates of $m$ marker positions (red shape in Fig. 2(b)) using Gauss’s area formula. We describe this as our characteristic area:
$A_{c}=\tfrac{1}{2}\left|\sum_{k=1}^{m}(x_{k}y_{k+1}-x_{k+1}y_{k})\right|,$ (1)
where $x_{m+1}=x_{1}$ and $y_{m+1}=y_{1}$.
We chose to analyze this characteristic area as the metric for the steerability of the robot under the various conditions tested, as opposed to either a traditional 3D workspace or a curvature measurement. A 3D workspace of the robot would plot all reachable robot tip positions, considering growth (one degree of freedom) and steering (two degrees of freedom). As a proxy for this 3D workspace measurement that allows us to examine trends as various parameters change, the characteristic area metric can be thought of as the area of a vertical slice through the 3D workspace of the robot, which represents the steerability of the robot at a given length. A curvature measurement (as we used in the supported planar steerability experiments described in Section IV) does not capture the robot’s behavior well when moving unsupported in 3D space, because the robot’s body deflects under the influence of gravity, not forming a constant curvature shape. Additionally, the robot’s range of motion is different in the horizontal and vertical directions, due to differing effects of gravity, so measuring the horizontal and vertical range of the positions reached at each constant length allows us to quantify that effect.
Fig. 3 shows the results of our self-supporting 3D steerability experiments. For each set of parameters tested (columns (a) through (d)), we plot the paths traced out by the robot tip during each trial (top row). We also plot the average characteristic area across three trials and its standard deviation (second row), along with the average and standard deviation for the horizontal range (the difference between the minimum and maximum $x$ values) (third row), the vertical range (the difference between the minimum and maximum $y$ values) (fourth row), and the initial vertical position (the $y$ value at the start of a trial, which begins when the robot reaches via point 1) for each trial (fifth row).
@@ III-C Tip Load
When varying the tip load, we maintained the following constant parameters: a chamber pressure of 6.89 kPa, which is high enough to support the robot while also being low enough for it to steer freely; a length of 0.61 m, which is sufficiently long to steer without causing the robot to excessively deflect under its own weight; and an inflated diameter of 8.1 cm, which is close to the size of previously deployed field robots [CoadRAM2020, der2021roboa, mcfarland2023collapse]. The robot performed its task under loads of 0 g, 20 g, 50 g, 100 g, 150 g, and 200 g (the maximum mass the robot could hold before it started approaching collapse behavior).
Steerability as a function of tip load is shown in Fig. 3(a). There is a clear decrease in characteristic area with increased tip load, with a total decrease of 89% occurring between a load of 0 g and a load of 200 g. Correspondingly, the horizontal range, vertical range, and initial vertical position decrease by 39%, 74%, and 0.38 m, respectively.
@@ III-D Chamber Pressure
When varying chamber pressure, we maintained the following constant parameters: a tip load of 50 g, an intermediate load that causes some deflection but not so little steerability as to prevent the analysis of other parameters; a length of 0.61 m; and a diameter of 8.1 cm. The robot performed its task at chamber pressures of 2.07 kPa (close to the minimum chamber pressure that supports the robot), 3.45 kPa, 5.52 kPa, 6.89 kPa, 8.27 kPa, and 10.34 kPa (close to the maximum chamber pressure it can hold without bursting).
Steerability as a function of chamber pressure is shown in Fig. 3(b). The characteristic area starts high at a low pressure and initially increases 14% with pressure, peaking at 5.52 kPa. Beyond this point, higher pressures lead to a 54% decrease in characteristic area compared to the initial value, as the robot is too stiff for the actuators to move it effectively. The horizontal range initially increases 3%, but ultimately decreases by 51% with increasing pressure, which is expected with the robot’s increasing stiffness. The vertical range generally increases around 26% with increasing pressure but experiences high variation. This is likely due to the fact that the robot is becoming stiffer and harder to move, but that same stiffness is also preventing deflection and therefore allowing the robot to make larger motions if it creates enough force to do so. Similarly, we see the initial vertical position increase 0.19 m over the pressure range, as a stiffer robot can start each trial with a higher tip position.
@@ III-E Length
When varying robot length, we maintained the following constant parameters: a tip load of 50 g; a chamber pressure of 6.89 kPa; and a diameter of 8.1 cm. The robot performed its task at lengths of 0.30 m, 0.46 m, 0.61 m, and 0.76 m.
Steerability as a function of robot length is shown in Fig. 3(c). Generally, the characteristic area increases with increasing length, with a 20% increase in characteristic area occurring from 0.30 m to 0.76 m. There is also an 80% increase in horizontal range, which makes sense, since a longer robot’s tip should travel farther for a given amount of curvature. However, the vertical range, which decreases by 30%, shows that this increase in length results in a heavier robot that struggles more to move against gravity. Eventually, the robot would become so long that it would collapse, reducing the characteristic area, horizontal range, and vertical range to nearly zero. We also see the initial vertical position dropping by 0.42 m, significantly limiting where in the vertical plane we can reach.
@@ III-F Diameter
When varying diameter, we maintained the following constant parameters: a tip load of 50 g; a chamber pressure of 6.89 kPa; and a length of 0.61 m. The robot performed its task at inflated diameters of 3.2 cm (smaller than the typical field exploration robot), 6.5 cm, 9.7 cm, and 12.9 cm (larger than the typical field exploration robot). It should be noted that the 9.7 cm diameter robot was oriented upside down for its trials, and therefore it formed a differently-oriented trajectory and had a correspondingly lower initial vertical position.
Steerability as a function of robot diameter is shown in Fig. 3(d). There is only a 4% change in characteristic area between the 6.5 cm diameter robot and the 12.9 cm diameter robot. The 3.2 cm robot collapsed under the given parameters. Thus, having a diameter large enough to avoid collapse is important, but how much larger is of minimal consequence. Similarly, the horizontal range, vertical range, and initial vertical position are fairly constant once the robot is at a sufficient diameter to not collapse, though there is a 17% decrease in horizontal range, a 16% increase in vertical range, and a 0.18 m increase in initial vertical position with increasing diameter, excluding the 3.2 cm robot.
@@ IV-A Robot Body Fabrication
In the exterior pouch design [CoadRAM2020], the main chamber is first fabricated as a single-layer thermoplastic polyurethane (TPU)-coated ripstop nylon tube. Three series pouch motor actuators, each with half the lay-flat width of the main chamber tube, and with pouches of equal length and width (experimentally determined to yield maximum contraction [kubler2024comparison]) are then constructed and taped outside the main chamber.
In the integrated pouch design [Ataka2020ModelBased], the main chamber and actuators are built together as a double-layer structure. A folded piece of TPU-coated fabric is heat-sealed to define both the body wall and internal pouch chambers before being joined lengthwise into a tube. In this construction, each series pouch motor actuator has a nominal width of two-thirds of the lay-flat diameter of the main chamber tube.